OpenStudy (anonymous):

How can I solve 4x + 8 over 2 = -10?

OpenStudy (anonymous):

Multiply the 2 out so 4x+8=-20 4x=-12 x=-3

OpenStudy (anonymous):

(4x+8) / 2 = -10 4x+ 8 = -20 4x = -12 x = -3

OpenStudy (anonymous):

Thanks dude :)

OpenStudy (anonymous):

4x+8 ------ =-10 2 multiply by 2 so u can get rid of denominators 4x+8=-20 4x=-12 x=-3

OpenStudy (anonymous):

can u give me a medal please